# Learning-Augmented Skip Lists

### v1

@@ 2.3 Zipfian Distribution
We next upper bound the entropy of a probability vector that satisfies a Zipfian distribution with parameter $s$.
Lemma 2.9.
@@
Since $f$ is a probability distribution on the support of size $[n]$, then by Theorem 2.8, we have that $H(f)\leq\log n$. Thus, it remains to consider the case where $s>1$. Since $z\leq 1$, we have
@@
By Lemma 2.5 and Lemma 2.9, we have the following statement about the performance of our learning-augmented skip list on a set of search queries that follows a Zipfian distribution.
Corollary 2.10.
With high probability, the expected search time on a set of queries with exponent $s$ is at most $\mathcal{O}\left(1\right)$ for $s>1$ and $\mathcal{O}\left(\log n\right)$ for $s\leq 1$.
3 Predictor Robustness
In this section, we modify our algorithm to be robust to arbitrarily inaccurate oracles, while still ensuring consistency when the oracle is accurate.
3.1 Noisy Robustness
Let $f$ be the true-scaled frequency vector so that for each $i\in[n]$, $f_{i}$ is the probability that a random query corresponds to $i$. Let $p$ be the predicted frequency vector, so that for each $i\in[n]$, $p_{i}$ is the predicted probability that a random query corresponds to $i$. For $\alpha,\beta\in(0,1)$, we call an oracle $(\alpha,\beta)$-noisy if for all $i\in[n]$, we have $p_{i}\geq\alpha\cdot f_{i}-\beta$.
Lemma 3.1.
Let $\alpha$ be constant and $\beta<\frac{1}{2n}$. A learning-augmented skip list with a set of $(\alpha,\beta)$-noisy predictions has performance that matches that of a learning-augmented learned with a perfect oracle, up to an additive constant.
@@ 3.1 Noisy Robustness
Suppose the total number of levels is at most $C+\log n$ for some constant $C>0$. Note that this occurs with a high probability for a learning-augmented skip list with a set of $(\alpha,\beta)$-noisy predictions. For each $i\in[n]$, let $f_{i}$ be the proportion of queries to item $i$ and let $p_{i}$ be the predicted proportion of queries to item $i$. By Lemma 2.5, the expected search time at most
@@
$p_{i}\geq\alpha\cdot f_{i}-\beta\geq\alpha\cdot f_{i}-\frac{\alpha}{4n}\geq\alpha\cdot\frac{\alpha}{2}\cdot f_{i}.$
@@
3.2 Consistency and Robustness
To ensure both robustness and consistency, we can provide the following adaptation. We first build a learning-augmented skip list ${\mathcal{L}}_{1}$ using the procedure in Algorithm 1. We then build a normal skip list for the items whose predicted values ${\mathcal{L}}_{2}$ are less than $\frac{1}{n}$, i.e., $p_{i}\leq\frac{1}{n}$. We then append ${\mathcal{L}}_{2}$ below the first level in ${\mathcal{L}}_{1}$.
Algorithm 2 Robust learning-augmented skip list
1:Predicted frequencies $p_{1},\ldots,p_{n}$ for each item in $[n]$
2:Learning-augmented skip list
3:Let ${\mathcal{L}}_{1}$ be the data structure from Algorithm 2
4:Let $S$ be the set of items in level 00 of ${\mathcal{L}}_{1}$
5:Build a skip list ${\mathcal{L}}_{2}$ for the items in $S$
6:Append ${\mathcal{L}}_{2}$ to the bottom of ${\mathcal{L}}_{1}$ by replacing level 00 from ${\mathcal{L}}_{1}$ with the top node of ${\mathcal{L}}_{2}$
We first show the consistency of our algorithm, i.e., good algorithmic performance when the oracle is accurate.
Lemma 3.2.
Algorithm 2 achieves a constant factor runtime of the optimal skip list when the predictions are correct.
@@ 3.2 Consistency and Robustness
We claim that when the predictions are correct, then Algorithm 2 will demonstrate the same performance as Algorithm 1. This is because an instantiation of Algorithm 1 creates the skip list ${\mathcal{L}}_{1}$, which is then inserted at the top of Algorithm 2. Because the predictions are correct, then the search process will terminate at the bottom level of ${\mathcal{L}}_{1}$, never needing to search through ${\mathcal{L}}_{2}$. By Lemma 2.5, we know that the expected search time through ${\mathcal{L}}_{1}$ is a constant factor of the runtime of the optimal skip list. Thus, the desired claim follows. ∎
We next show the robustness of our algorithm, i.e., standard algorithmic performance when the oracle is inaccurate.
Lemma 3.3.
Algorithm 2 achieves a constant factor runtime of an oblivious skip list construction when the predictions are incorrect.
@@
We first note that ${\mathcal{L}}_{1}$ has height $\mathcal{O}\left(\log n\right)$. Moreover, an oblivious skip list construction has expected search time $\mathcal{O}\left(\log n\right)$. Since ${\mathcal{L}}_{2}$ is a standard skip list construction, then in expectation, the total search time is $\mathcal{O}\left(\log n\right)$, even when the predictions are incorrect. Therefore, the desired statement holds. ∎
4 Empirical Evaluations
@@
In this section, we perform empirical evaluations comparing the performance of our learning-augmented skip list to that of traditional skip lists, on both synthetic and real-world datasets. Firstly, we compare the performance of traditional skip lists with our learning-augmented skip lists on synthetically generated data following Zipfian distributions. The proposed learning augmented skip lists are evaluated empirically with both synthetic datasets and real-world internet flow datasets from the Center for Applied Internet Data Analysis (CAIDA) and AOL. In the synthetic datasets, a diverse range of element distributions, which are characterized by the skewness of the datasets, are evaluated to assess the effectiveness of the learning augmentation. In the CAIDA datasets, the $\alpha$ factor is calculated to reflect the skewness of the data distribution.
@@
4.1 Synthetic Datasets
In the synthetic datasets, both the vanilla and augmented skip lists are tested against different element counts and $\alpha$ values. In terms of the distribution of the synthetic datasets, the uniform distribution and a Zipfian distribution of $\alpha$ between 1.01 and 2 with query counts up to 4 million are evaluated. It is worth noting that the number of unique element queries could vary for the same query count at different $\alpha$ values in the Zipfian distribution, which may affect the insertion time.
Table 1: Speed up factor of augmented skip list over vanilla skip list under different synthetic distributions
@@ 4.1 Synthetic Datasets
Table 1 shows the speed-up factor, defined as the time taken by the augmented skip list over the vanilla skip list for the same query stream. We can observe a progressive improvement in the performance of our augmented skip lists as the dataset skewness increases. It also suggests that our augmented skip list will perform at least as good as the traditional skip list and will outperform a traditional skip list by a factor of up to 7 times depending on the skewness of the datasets.
@@
The query time of augmented skip lists is also reduced greatly compared to the vanilla skip lists as shown in Figure 2.
4.2 CAIDA Dataset
In the CAIDA datasets [CAI16], the receiver IP addresses from one minute of the internet flow data are extracted for testing, which contains around 650k unique IP addresses of the 30 million queries.
Given that the log-log plot of the frequency of all nodes in the CAIDA datasets follows approximately a straight line in Figure 3, the CAIDA datasets can be approximately characterized by an $\alpha$ factor of 1.37.
The insertion time is similar between vanilla and augmented skip lists, while the query time is greatly reduced according to Figure 4.
It is worth noting that the results presented above assume that the frequency of all items in the query stream is accurate, which means the probability vector that is used to build the skip list matches exactly the query stream. However, this can be challenging to achieve in dynamic datasets where the probability vector may change over time. We hereby demonstrate that our proposed algorithm still manages to outperform the vanilla skip list even when temporal change exists in the probability vector. We demonstrate this by comparing the query time for the same set of query elements with different probability vectors being used to guide the building of the structure. The results are presented in Figure 4(a).
In Figure 4(a), color bars indicate the different skip list structures being compared, namely the vanilla skip list, the skip list augmented by noisy probability vectors, and the skip list augmented by perfect learning. In the vanilla skip list, as explained earlier, all elements are promoted to higher levels with equal chance. For the skip list augmented by a noisy probability vector, the probability vector of elements during a period of T1 is used as the reference probability vector. The skip list being augmented by this probability vector has its own set of elements to be organized into the target skip list. Suppose the historic data from T1 contains a set of elements S1, and some future query stream contains a set of elements S2. For each element in our target set S2, if the element is present in S1, then the occurrence probability of this element from S1 will be used to build S2; otherwise, if the element has not shown up during T1 (i.e., in S1), then we assume its probability to be 0. After this, the probability vector is normalized to sum to 1, resulting in a predicted probability vector to be used to build a skip list based on the historic element frequency. The dict-type object which maps the element value to a probability value is also referred to as an oracle. Since there is temporal changes in the frequency of elements being queried, the predicted probability vector will show a discrepancy with the true probability vector. We show that even when the prediction is not perfect, the augmented skip list still performs better than a conventional skip list. The skip list with perfect learning means that the probability vector which is used to augment the skip list matches exactly the occurrence probability of queried items in the query stream.
The CAIDA datasets used in this study contain 12 minutes of internet flow data, which totals around 444 million queries. The indices on the x-axis in Figure 4(a) means:
Figure 4(a) shows that the skip list with perfect learning shows the best performance, while the skip list augmented with noisy learning performs very close to the perfect learning scenario. It is also noted the closer the test data is to the reference data chronologically, the closer the noisy-augmented skip list will perform to the perfect learning skip list.
4.3 AOL Dataset
@@
Fig. 1: Insertion time for synthetic datasets with a uniform distribution and under different $\alpha$ values of the Zipfian distribution for both vanilla and augmented skip lists. This figure illustrates the insertion time on the synthetic data for both the uniform distribution and the Zipfian distribution at different $\alpha$ values. Generally, higher skewness of the datasets results in less insertion time when using the augmented structure. The decrease in insertion time is proportional to the increase in the $\alpha$ value, as a higher $\alpha$ value leads to a reduction in the number of unique nodes, as illustrated in Table 2.
@@
Fig. 2: Query time for synthetic datasets with a uniform distribution and under different $\alpha$ values of the Zipfian distribution for both vanilla and augmented skip lists. This figure compares the query time of the vanilla and augmented skip lists for both uniform distribution and Zipfian distribution at various $\alpha$ values. Similar to insertion, the query time is significantly reduced under different query sizes with the implemented augmentation. The performance enhancement is especially pronounced for the high skewness of the dataset.
(a) CAIDA data distribution
(b) Zipfian fit ($\alpha=1.37$)
Fig. 3: CAIDA datasets distribution characterization. This figure illustrates the derivation of the $\alpha$ value, determined to be 1.37, for the CAIDA dataset. The nearly straight-fitted curve supports the reasonability of characterizing the CAIDA datasets as a Zipfian distribution.
(a) Insert time on CAIDA
(b) Query time on CAIDA
Fig. 4: Insertion and query time on CAIDA for vanilla and augmented skip lists. This figure compares the insertion and query times under varying numbers of top frequently accessed unique IPs between vanilla and augmented implementations. The horizontal axis in subfigures a and b of Figure 4 depicts the same scheme of IP selection, represented in two different ways. For example, the top 29.9 million queries contain 665210 unique IPs, and the top 29.5 million queries comprise 296384 unique IPs, and so forth. The query time is almost halved when using the augmented skip lists at different query sizes.
(a) Robustness test on CAIDA datasets
(b) Oracle credibility on CAIDA datasets
Fig. 5: CAIDA datasets distribution characterization. This figure demonstrates the robustness of our algorithm even in the presence of a non-perfect oracle. The augmented skip list exhibits performance closely comparable to the perfect oracle, and for CAIDA datasets, minimal temporal changes in the probability vector result in performance closely approaching that of the perfectly augmented skip list. In subfigure b, the labels on the axis indicate the time stamp that the internet trace data is collected. For example, 130100 means the collection starts at 13:01:00 and lasts for 1 minute. The subfigure b reads 94% of the queries between timestamps 131000 and 125910 are the same. This indicates that the CAIDA data is relatively stable, which explains the robustness of our oracle.

### v2

@@ B.3 Zipfian Distribution
We can then upper bound the entropy of a probability vector that satisfies a Zipfian distribution with parameter $s$. See 2.3
@@
Since $f$ is a probability distribution on the support of size $[n]$, then by Theorem B.8, we have that $H(f)\leq\log n$. Thus, it remains to consider the case where $s>1$. Since $z\leq 1$, we have
@@
By Theorem 2.1 and Lemma 2.3, we have the following statement about the performance of our learning-augmented skip list on a set of search queries that follows a Zipfian distribution. See 2.4
B.4 Noisy Robustness
In this section, we show that our learning-augmented skip list construction is robust to somewhat inaccurate oracles. Let $f$ be the true-scaled frequency vector so that for each $i\in[n]$, $f_{i}$ is the probability that a random query corresponds to $i$. Let $p$ be the predicted frequency vector, so that for each $i\in[n]$, $p_{i}$ is the predicted probability that a random query corresponds to $i$. For $\alpha,\beta\in(0,1)$, we call an oracle $(\alpha,\beta)$-noisy if for all $i\in[n]$, we have $p_{i}\geq\alpha\cdot f_{i}-\beta$. See 2.5
@@ B.4 Noisy Robustness
Suppose the total number of levels is at most $C+\log n$ for some constant $C>0$. Note that this occurs with a high probability for a learning-augmented skip list with a set of $(\alpha,\beta)$-noisy predictions. For each $i\in[n]$, let $f_{i}$ be the proportion of queries to item $i$ and let $p_{i}$ be the predicted proportion of queries to item $i$. By Lemma B.5, the expected search time at most
@@
$p_{i}\geq\alpha\cdot f_{i}-\beta\geq\alpha\cdot f_{i}-\frac{\alpha}{4n}\geq\frac{\alpha}{2}\cdot f_{i}.$
@@
Appendix C Missing Proofs from Section 3
First, we will show that the expected depth of a given query depends on the probability of that query, and that high frequency queries must be found close to the root of our tree.
Lemma C.1.
Suppose $[\Delta]^{d}$ is the space of possible input points and queries. Let $N=\Delta^{d}$ and $p_{i}$ be the probability that a random query is made to $i\in[N]$, given the natural mapping between $[N]$ and $[\Delta]^{d}$. Then the level at which $i$ resides in the tree is at most $\mathcal{O}\left(\log\frac{1}{p_{i}}\right)$.
@@ Appendix C Missing Proofs from Section 3
First, consider only the high-frequency query points and data points for which we base our construction off.
In constructing the learning-augmented KD tree, we balance the contents of the children nodes such that a query to that node has a probability of $\frac{1}{2}$ of belonging to each of the children. Therefore, at a depth of $d$, the probability of belonging to either child is $\frac{1}{2^{d}}$. In particular, a query $i$ with probability $p_{i}$ at a depth $d$ must satisfy $p_{i}>\frac{1}{2^{d}}$. Thus, we have that the depth of $i$ is $\mathcal{O}\left(\log\frac{1}{p_{i}}\right)$, as desired.
Since the lowest probability of a high-frequency data point is $\frac{1}{n^{2}}$, this tree must have a depth of at most $2\log n$.
Now, consider a low-frequency data point, which we add to the bottom of the tree. By construction, the learned point of our tree has depth at most $2\log n$. Then, when inserting the additional data points as a balanced KD tree, we can accumulate at most an additional depth of $\log n$. Note, $p<\frac{1}{n^{2}}$ implies $\log n<\log\frac{1}{p}$. Thus, this low-frequency data point will have a depth of at most $3\log n=\mathcal{O}\left(\log\frac{1}{p}\right)$, as desired.
Similarly, if $i$ is not a data point and is low frequency, we achieve the same bound of $\mathcal{O}\left(\log\frac{1}{p}\right)$. In this case, we simply terminate at a leaf node and determine that the desired query is not in the dataset.
In summary, any query which has high frequency can be found in $\mathcal{O}\left(\log\frac{1}{p}\right)$ time. Low-frequency data points can similarly be found in $\mathcal{O}\left(\log\frac{1}{p}\right)$ time, and low frequency queries can be determined to not exist in $\mathcal{O}\left(\log\frac{1}{p}\right)$ time. ∎
Lemma C.2.
Suppose $[\Delta]^{d}$ is the space of possible input points and queries. Let $N=\Delta^{d}$ and $p_{i}$ be the probability that a random query is made to $i\in[N]$, given the natural mapping between $[N]$ and $[\Delta]^{d}$. Then the level at which $i$ resides in the tree is at most $\mathcal{O}\left(\log n\right)$.
@@
This follows directly from the analysis in Llemma C.1. ∎
Now, we have demonstrated the the depth of a given query point $i$ is bounded by both $\mathcal{O}\left(\log n\right)$ and $\mathcal{O}\left(\log\frac{1}{p_{i}}\right)$. Using this fact, we will now show that the expected query time of our algorithm is bounded by both the entropy of the dataset $H(p)$ in addition to $\log n$.
We now analyze the performance of our learning-augmented KD tree. See 3.1
Proof.
Following C.1, the points $i$ in $[\Delta]^{d}$ with non-negligible probability $p_{i}\geq\frac{1}{n^{2}}$ are guaranteed to exist in the learning-augmented KD tree $\mathcal{T}$ with depth at most $\log\frac{1}{p_{i}}$. For points in the dataset $[n]$ with negligible probability, they exist in the tree and have depth in $\mathcal{O}\left(\log n\right)$. For all other points not contained in the KD tree, the query will terminate at a depth of $\mathcal{O}\left(\log n\right)$.
For any point $i$ in $[\Delta]^{d}$, the depth that a query to $i$ will terminate in the tree is
$\operatorname*{Depth}(i)=\mathcal{O}\left(\min\left(\log\frac{1}{p_{i}},\log n\right)\right).$ (1)
Then, the expected search time $T$ is the expected depth of a given point,
$\mathbb{E}\left[T\right]=\sum_{i\in[\Delta]^{d}}p_{i}\cdot\operatorname*{Depth}(i)=\sum_{i\in[\Delta]^{d}}p_{i}\cdot\mathcal{O}\left(\min\left(\log\frac{1}{p_{i}},\log n\right)\right)=\mathcal{O}\left(\min(H(p),\log n)\right).$ (2)
Thus, we have shown that the expected query time of our algorithm is bounded by the entropy of the dataset. In particular, when the dataset has a highly skew distribution, $H(p)$ can be far less than $\log n$.
C.1 Near-Optimality
Near-optimality of our learning-augmented KD trees uses a similar argument to the proof of the near-optimality of our learning-augmneted skip lists. In particular, we again utilize Shannon’s source coding theorem from Theorem B.7. We then have the following: See 3.2
Proof.
In a learning-augmented KD tree, the search path to an element $i$ can be encoded as a $0-1$ codeword, with entries indicating whether the lower or upper branch is taken at each node traversal. Moreover, the length of this codeword in the symbol-by-symbol encoding corresponds to the depth of element $i$. Then, by Theorem B.7 and the optimality of Huffman codes in symbol-by-symbol encodings, we have that $\mathbb{E}\left[D(X)\right]\geq H(f)$, as desired. ∎
C.2 Noisy Robustness
In this section, we analyze the performance of the learning-augmented KD tree under noisy data conditions.
Previously, we analyzed the performance of the learning-augmented KD-tree assuming access to a perfect prediction oracle.
Now we analyze the performance with a noisy oracle. That is, for each $i\in[\Delta]^{d}$ there is a true-scaled frequency $f_{i}$ that the point will be queried and that $p_{i}$ is a prediction made by the noisy oracle.
First, we analyze the multiplicative robustness of the algorithm. In this case, the oracle predicts $f_{i}$ up to some multiplicative constant $\alpha\in\mathbb{R}^{+}$ such that $f_{i}=\alpha\,p_{i}$.
Lemma C.3.
Suppose $[\Delta]^{d}$ is the space of possible input points and queries. Let $N=\Delta^{d}$ and $p_{i}$ be the probability that a random query is made to $i\in[N]$ during tree construction. Suppose during runtime that the true probability of querying $i$ is $f_{i}=\alpha\,p_{i}$ for some $\alpha\in\mathbb{R}^{+}$. Then the level at which $i$ resides in the tree is at most $\mathcal{O}\left(\log\frac{1}{f_{i}}+\log\frac{1}{\alpha}\right)$.
Proof.
If $\alpha\leq 1$, this is immediate. If this is the case, in construction we expected $i$ to be queried more often than it actually is, so our construction placed the point $i$ higher in the tree than is necessary. Thus, the depth is at most the previously shown $\frac{1}{p_{i}}\leq\frac{1}{f_{i}}$.
Now, suppose $\alpha>1$. In this case, we must have placed $i$ deeper in the tree than we should have, as our construction frequency is less than the true query frequency. Then, as in C.1, the depth of $i$ is $\mathcal{O}\left(\log\frac{1}{p_{i}}\right).$ Now, we have that
$\mathcal{O}\left(\log\frac{1}{p_{i}}\right)=\mathcal{O}\left(\log\frac{\alpha}{f_{i}}\right)=\mathcal{O}\left(\log\frac{1}{f_{i}}+\log\alpha\right).$ (3)
Having shown multiplicative robustness of the learning-augmented KD tree we next analyze the additive-multiplicative robustness of the method. An oracle is $(\alpha,\beta)$-noisy if the prediction satisfies $p_{i}\geq\alpha\,f_{i}-\beta$ for constants $\alpha,\beta\in(0,1)$.
See 3.4
Proof.
From Lemma C.1 we have that the depth of a point $i$ in the learning-augmented KD tree is $\mathcal{O}\left(\frac{1}{p_{i}}\right)$ where $p_{i}$ is the predicted frequency of $i$.
Then, with an $(\alpha,\beta)$-noisy oracle we have that the prediction is bounded from below by $p_{i}\geq\alpha\,f_{i}-\beta$. As $i$ is a point in the tree, we can assume that the true-scaled frequency $f_{i}$ has a lower-bound of $1/n^{2}$.
By choosing $\beta\leq\alpha/n^{2}$ we ensure that the predicted $p_{i}$ is always nonnegative. Then, let $\beta=\alpha/2n^{2}$
$p_{i}\geq\alpha\,f_{i}-\beta\geq\alpha\,f_{i}-\frac{\alpha}{2n^{2}}\geq\frac{\alpha}{2}f_{i}$ (4)
Then, applying Lemma C.1 again
$\mathcal{O}\left(\log\frac{1}{p_{i}}\right)=\mathcal{O}\left(\log\frac{2}{\alpha f_{i}}\right)=\mathcal{O}\left(\log\frac{1}{f_{i}}+\log\frac{1}{\alpha}\right).$ (5)
Appendix D Additional Empirical Evaluations
D.1 Skip Lists
@@
In this section, we perform empirical evaluations comparing the performance of our learning-augmented skip list to that of traditional skip lists, on both synthetic and real-world datasets. Firstly, we compare the performance of traditional skip lists with our learning-augmented skip lists on synthetically generated data following Zipfian distributions. The proposed learning-augmented skip lists are evaluated empirically with both synthetic datasets and real-world internet flow datasets from the Center for Applied Internet Data Analysis (CAIDA) and AOL. In the synthetic datasets, a diverse range of element distributions, which are characterized by the skewness of the datasets, are evaluated to assess the effectiveness of the learning augmentation. In the CAIDA datasets, the $\alpha$ factor is calculated to reflect the skewness of the data distribution.
@@
D.1.1 Synthetic Datasets
In the synthetic datasets, both the classic and augmented skip lists are tested against different element counts and $\alpha$ values. In terms of the distribution of the synthetic datasets, the uniform distribution and a Zipfian distribution of $\alpha$ between 1.01 and 2 with query counts up to 4 million are evaluated. It is worth noting that the number of unique element queries could vary for the same query count at different $\alpha$ values in the Zipfian distribution, which may affect the insertion time.
Table 1: Speed up factor of augmented skip list over classic skip list under different synthetic distributions
@@ D.1.1 Synthetic Datasets
Table 1 shows the speed-up factor, defined as the time taken by the augmented skip list over the classic skip list for the same query stream. We can observe a progressive improvement in the performance of our augmented skip lists as the dataset skewness increases. It also suggests that our augmented skip list will perform at least as good as the traditional skip list and will outperform a traditional skip list by a factor of up to 7 times depending on the skewness of the datasets.
@@
The query time of augmented skip lists is also reduced greatly compared to the classic skip lists as shown in Figure 7.
@@
Figure 6: Insertion time for synthetic datasets with a uniform distribution and under different $\alpha$ values of the Zipfian distribution for both classic and augmented skip lists. This figure illustrates the insertion time on the synthetic data for both the uniform distribution and the Zipfian distribution at different $\alpha$ values. Generally, higher skewness of the datasets results in less insertion time when using the augmented structure. The decrease in insertion time is proportional to the increase in the $\alpha$ value, as a higher $\alpha$ value leads to a reduction in the number of unique nodes, as illustrated in Table 2.
@@
Figure 7: Query time for synthetic datasets with a uniform distribution and under different $\alpha$ values of the Zipfian distribution for both classic and augmented skip lists. This figure compares the query time of the classic and augmented skip lists for both uniform distribution and Zipfian distribution at various $\alpha$ values. Similar to insertion, the query time is significantly reduced under different query sizes with the implemented augmentation. The performance enhancement is especially pronounced for the high skewness of the dataset.
In addition, we conduct experiments to compare the performance of standard binary search trees and standard skip lists. In particular, we generate datasets of size $n\in\{5000,10000,15000,20000,25000,30000,35000,40000,45000,50000,\}$. For each fixed value of $n$, each element of the dataset is generated uniformly at random in $[2n]$, i.e., uniformly at random from $\{1,2,\ldots,2n\}$. Because the dataset is generated uniformly at random, then learning-augmented data structures will perform similar to oblivious data structures. We measure the construction time of the data structures, based on the input dataset. Our results demonstrate that as expected, skip lists perform significantly better than balanced binary search trees across all values of $n$, due to the latter’s necessity of constantly rebalancing the data structure. In fact, skip lists performed almost 4×4$\times$4 × better than BSTs in some cases, e.g., $n=20000$. We illustrate our results in Figure 8.
Figure 8: BST vs skip list construction times
D.1.2 AOL Dataset